Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎


Crew 100B Log Book for March 8, 2011

Photos of the day

Commander's Check-In Report

Balwant Rai Reporting

Physical Status: Excellent

Experiment Status: All the medical and simulated micro-gravity , our habitability began today with unique results.

Time Departed/Returned from EVA: One EVA performed

The meals today were OK since it was a non-cooking day.

EVA’s: We completed one successful EVA.

Number Of reports: 3 science reports, 1 engineering report, 1 journalist report, 1 EVA report and 1 chef report.

Plan for tomorrow: will initiate Human machine interaction experiment again on all crew members. Interpretation of the Sun by means of the radio telescope will also commence. The study of the rover’s maneuverability as well as the potential use of the drone, weather pending, will take place. Soil samples will yet again be collected during three EVA’s tomorrow. Tomorrow will be a cooking day.

Commander's Report

Balwant Rai Reporting

Today we started the day with simulated microgravity and medical experiments and our habitability experiment.

The crew’s medical health will continue to be monitored by our health and safety officer for the remainder of our stay here on Mars. Breakfast was satisfying. Our crew had their saliva samples taken (via Versi Sal 1 from Oasis Diagnostics), vital parameters monitored (via the Zephyr Bioharness), and their heart rate variability measured (via NERV Express 4.2 software). I also performed the cognitive study (via CogState Research software) on all crew members, as well.

Our crew scientist continued to work through problems with starting observations of the sun via the radio telescope. One EVA was performed to sites of interest. Today is a non-cooking day. Our astronomer continued her experiments well. Today was Simulated micro-gravity day.

Human machine, simulated microgravity experiments, Medical, habitability and astrobiological experiments will be taken into account, also, the rover experiment and the radio telescope experiments will be performed.

Engineering Report

Matthieu Ansart Reporting


· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good

Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): full

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.34

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 2 containers full



· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):

Water Status:

· Outside Potable Water Tank Level (inches from bottom): 6

· Trailer Potable Water Tank Level (inches from bottom): Empty

· Hab Potable Water Tank Level (inches from bottom): 13

· Potable Water Meter Reading: 432168

Notes/Comments : Our water level is very low.


· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green plants and snails

· Condition of plants in Tank 3: green plants

· Greenhab notes/ comments:


· Telescope Used? (Yes/no) No

Observatory Notes:



· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:

ATV 1 (Opportunity):

· Used: No

· Oil Checked: no

· Fuel Consumed:no

· Tire Status: not checked

· Comments:

ATV 2 (Viking I):

· Used: No

· Oil Checked: no

· Fuel Consumed: no

· Tire Status: not checked

· Comments:

ATV 3 (Viking II):

· Used: No

· Oil Checked: no

· Fuel Consumed: None

· Tire Status: not checked

· Comments:

Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 73

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 59

Computers/Networking Infrastructure:

Notes/comments on internet/computers?:

General Engineering

1. Notes/Comments:

Science Report (Bio-Medical)

Balwant Rai and Jasdeep Kaur Reporting

JBR group study of Simulated Microgravity Bio-medical Experiments: MDRS Crew 100B ILEWG EuroMoonMars Crew

Marsonauts will stumble upon both physiological and psychological extremes during the journey while on the Mars terrain and the journey back to Earth. Considering these factors, short duration analogue studies, such as those being accomplished at the MDRS, Utah, USA, propose an opportunity to study mission operations and human factors in a simulated environment and contribute to plan missions to explore the Moon and Mars (MDRS Crew 100B ILEWG EuroMoonMars). The MDRS Crew 100B ILEWG EuroMoonMars, performed 6 hour HDT ( simulated microgravity condition) and provided a unique insight into physiology issues for space exploration. In this study, Salivary biomarkers ( bone loss and formation stress, immunological and inflammatory markers: saliva samples were taken by via Versi Sal 1 from Oasis Diagnostics), questionnaire, cognitive activity (CogState Research software), vital parameters were monitored (via the Zephyr Bioharness), and heart rate variability was measured (via NERV Express 4.2 software). These factors were taken into account and analyzed by subjective and objective means during 100BILEWG EuroMoonMars and preliminary results of all were summarized. We concluded from this study results that the cumulative, long-duration burdens associated with an astronaut’s health-capacity and their ability and tolerance to adjust to all mission conditions imposes considerable concerns that have to be further investigated before human Mars exploration advances.

Science Report (Medical)

Jasdeep Kaur Reporting

Time: 0930-Blood pressure and pulse rate monitored by Zephyr technology and by manually prior to and after IVA experiments.

Crew Members: Science Team: 6 crew members

Location: Hab

Objective: To examine the affects of different IVA activities on blood pressure and pulse rate. Observation: All crew members recorded the blood pressure and pulse rate early in the morning and then different medical and other habitability experiments were performed. The blood pressure was every-time recorded after the completion of experiments.

Results:There is a very little difference in blood pressures examined after different IVA activities in all crew members, this is due to adaptability.

Lessons learned: Vital parameters should be taken into account very carefully.

Science Report (Astronomy)

Crystal Latham Reporting

Observations with the radio telescope have been going fairly well. Two days ago there was a recorded solar burst, however, it was prior to sunrise here on Mars and it was not measured. Solar bursts and any activity from the Sun is very sporadic and it cannot be predicted like radio activity from Jupiter. It is a waiting game. Since the Sun has been getting more active it is getting more and more likely solar bursts will occur that can be recorded here at MDRS.

Since, there have been no recorded solar bursts I have started using the time I am recording radio frequencies with the telescope to determine what types of interference affect the telescope and why. MDRS is in a very secluded location in the Utah desert, however, that does not mean radio interference can’t reach us on analog Mars.  I am finding there are a number of sources of interference both inside and outside the Hab. Some of the interference I am recording is eluding classification, for now, while others are quite obvious and controllable.

EVA 32 Report

Rachel Dompnier Reporting

EVA #32, 4:30 to 5:00 pm

Crew Members: Matthieu Ansart, Quentin Bourgès, Rachel Dompnier

Site Location: in the vicinity of the HAB

Transit Mode: pedestrian EVA

Objectives: Take samples.

Results: We managed to take samples in a former river bed.

Lesson learned: We know how to use the GPS very well.

Chef's Report

Jasdeep Kaur Reporting

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?


2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Texas Barbecue chicken with beans

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)


4) Recipes for anything more complicated than following package directions exactly, or rehydrating.


5) Lessons learned (if any)


6) Comments/questions for Kim and Jean, the food study investigators


Journalist Report

Quentin Bourges Reporting

This morning, the 100B Crew woke up very enthusiastic and ready to attack a big working day. But before going to experiments, the Crew ate a very nice breakfast in order to get energy for the whole day.

We had many experiments to conduct. We performed a microgravity experiment which was supervised by Balwant, our Commander. We also did a pedestrian EVA to get some samples to analyze. Crystal, our Chief Scientist, managed to get some strange results with her radio telescope experiment. A number of interferences are being created, but we are unsure what they are from. Hence, she has planned to lead an investigation around the Hab tomorrow to figure out what exactly is causing these interferences. Unfortunately for the cadets, the Martian weather was cooler and windier than it has been for these three previous days, so they didn’t risk taking the UAV outside. The Crew has also nearly completed the habitability experiment.

There were enough leftovers cooked by our Chief Chef, Jasdeep, from yesterday to have a good lunch. As today was a non-cooking day, it was also a Strumstick day as we found this little music instrument left by Crew24 behind the stereo. Our executive officer had the opportunity to show his talents for string instruments by playing “Amazing Grace”. Today was a far more relaxing than many days here on Mars.